OpenStudy (anonymous):

Solve -6(x + 5) ≤ -x + 11. Write your solution using set builder notation

OpenStudy (anonymous):

Distribute the -6 and solve for x. -6x-30<=-x+11 -5x<=41 x>=-41/5 \[\left\{ x \in \mathbb{R}|x \ge-41/5 \right\}\]. I believe that is the right set notation.

OpenStudy (anonymous):

how do i make those symbols?

OpenStudy (anonymous):

Click "Equation" and then click through until you find them.

OpenStudy (anonymous):

would it be incorrect if i was to jus {x > -41/5} -

OpenStudy (anonymous):

equation? from where

OpenStudy (anonymous):

Well, the beginning says, x is any real number then the rest, such that x>-41/5. It is important. Depends on your teacher. But I would have it.

OpenStudy (anonymous):

ok